OpenStudy (anonymous):

what is the ordered pair, solve using matrix 7x+3y=58 6x+6y=72

OpenStudy (binary3i):

\[\left[\begin{matrix}7 & 3 \\ 6 & 6\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right) = \left(\begin{matrix}58 \\ 72\end{matrix}\right)\] x=4 y=8